OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem? 3n2 + 5 = 116 3n2 + 3n + 3 = 116 3n2 + 6n + 5 = 116

OpenStudy (anonymous):

How do you think would you solve this?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Well, think about it. Let's say the first number is \(n\), then the next is \(n + 1\), and the last one will be \(n+2\)

OpenStudy (anonymous):

sooooo n+n+1+n+2=116 or no???

OpenStudy (triciaal):

The sum of the squares

OpenStudy (triciaal):

n^2 + (n + 1)^2 + (n+ 2)^2 = 116

OpenStudy (anonymous):

ok sooo divide the equation by two right

OpenStudy (triciaal):

no

OpenStudy (triciaal):

n^2 + (n + 1)^2 + (n+ 2)^2 = 116

OpenStudy (triciaal):

expand the parenthesis group like terms you are just adding the sum

OpenStudy (triciaal):

3 n^2 + (2 + 4) n + (1 + 4) = 116 3 n^2 + 6 n + 5 = 116 Choice D

OpenStudy (anonymous):

ohh ok thanks

OpenStudy (triciaal):

you are welcome What are the numbers? do you need ?